OpenStudy (anonymous):

Three multiplied by the sum of 4 and a number is the same as 18 more than the number. Find the number. If n is "the number," which equation could be used to solve for the number? 3n + 4 = n + 18 4(3 + n) = n + 18 3(4 + n) = n + 18

OpenStudy (anonymous):

That sounds like the way to solve for n. 12 +3n = n +18 3n -n = 18-12 2n = 6 n =3 3 would be the number, but do you understand how to translate the given written form to get 3(4+n) = n + 18

OpenStudy (anonymous):

The sum of 4 and a number is 4 + n Three times that is 3(4 + n) 18 more than a number is n + 18 3(4 + n) = n + 18